2.Evaluate ∫cos(6x+4)dx
a.sin(6x+4)
--------- +c
6

b.cos(6x+4)
--------- +c
6
c.tan(6x+4)
--------- +c
6
d.sec(6x+4)
--------- +c
6
1
Expert's answer
2019-04-16T10:05:54-0400

Solution. let y=6x+4 get dy=6dx. We make a substitution and get



"\int cos(6x+4)dx=\frac {1} {6} \int cos(y)dy"

Using a table of integrals


"\int cos(x)dx=sin(x)+C"

where C is constant. Therefore


"\frac {1} {6} \int cos(y)dy=\frac {1} {6} sin(y)+C"

Returning to the substitution we get


"\frac {1} {6} sin(y)+C=\frac {1} {6} sin(6x+4)+C"


where C is constant.

Answer. a.

"\frac {1} {6} sin(6x+4) + C"

where C is constant.